OpenStudy (anonymous):

The midpoint of segment XY is Z(6, -3). The coordinates of one endpoint are X(-1, 8). Find the coordinates of endpoint Y. MEDALS!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (solomonzelman):

Plug in the numbers into mid point formula.

OpenStudy (anonymous):

UMM but would thant give me the mid point of yz i need to find y

OpenStudy (anonymous):



OpenStudy (solomonzelman):

\[\huge\color{red}{ (\frac{x_1+x_2}{2} ~~~,~~~\frac{y_1+y_2}{2})}\]

OpenStudy (anonymous):

i know th formula but how do i get y

OpenStudy (solomonzelman):

\[\huge\color{red}{ (\frac{-1+x_2}{2} ~~~,~~~\frac{8+y_2}{2})=(6,-3)}\]

OpenStudy (solomonzelman):

\[\huge\color{red}{ \frac{-1+x_2}{2}=6 ~~~~~~\frac{8+y_2}{2}=-3}\]

OpenStudy (anonymous):

ohhhh 14, -14

OpenStudy (anonymous):

right

OpenStudy (solomonzelman):

Oh, I was in the abthrrr \[\color{red}{ \frac{-1+x_2}{2}=6 ~~~~~~\frac{8+y_2}{2}=-3}\]\[\color{red}{ -1+x_2=12 ~~~~~~8+y_2=-6 }\]\[\color{red}{ x_2=13 ~~~~~~y_2=-14 }\]

OpenStudy (anonymous):

i was close, thank you

OpenStudy (solomonzelman):

Yeah you were, good job, yw!